From GlueXWiki
Jump to: navigation, search

June 11, 2015 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. CDC analysis update (Beni, Simon)
    1. Alignment: Fall'14: [2] Spring'15:[3]
  2. FDC analysis update (Lubomir)
  3. Electronics
  4. Engineering
    • Power outage preparations
  5. Other


Participants: Curtis (CMU), Sean (NU), Eugene, Luke, Dave, Chris, Nick, Simon, Beni, Fernando, and Lubomir (JLab).

CDC analysis update

- Simon showed new alignment results and compared them to the results for the fall 2014 run. For the new alignment he used external (BCAL) tracking, for the old - tracking from the CDC itself. One can see that in the new results the offsets are smaller - few hundreds microns, while in the old results there are groups of straws with offsets of ~600 microns.

- Curtis has seen recent results from Mike using Millepede package: the offsets are randomly distributed, not in groups as it is in the Simon's results. Will discuss this when Mike is back.

FDC update

- Lubomir looked at the ratio of the charge on the upper and down cathodes from the last run data. Plots above show the ratio for the first and second package - certainly there are systematic variations as a function of the wire position. Similar variations were observed for the residuals of the cathode reconstructed wire positions, but for the up/down ratio the variations are different in magnitude for the different packages. We discussed whether it may come from the wires having variation in their up/down positioning (defined by the inner side of the PC boards where the wires are glued), or these are some effects of the cathodes.

- It has been decided to continue the studies of the TR with the spare FDC package, this time it will be installed in the Hall on the platform at the pair spectrometer left arm (looking downstream). The shielding behind the PS counters will be modified to have an opening for the spare package placed behind it and additional shielding will be placed right behind the package. Fernando asked about extra electronics/cables needed for this; most of the stuff exists in the present set-up in the tagger.


- Nick: in the past weeks a CDC HV module was replaced, due to one channel not seeing the right hardware limit. However, the new HV module showed a weired behavior - voltage not stable on all the channels - so the module was replaced, as well.


- In preparations for the power outage Beni reduced the gas flow of the CDC by 50% and of the FDC by 40%. Luke verified that the chambers are still bubbling. According to the tests last week done by Mark Stevens and Dave, this will be enough for the gas system with the new UPS to survive the 4h power outage. For the long power outage generators will be used. All the HV and LV are off and also Beni turned off all the VXS and HV crates.